# Time-frequency Entangled Photon Mediated CCZ Gate

### v1

@@ I Introduction
Despite significant advances in superconducting quantum computing over the past few decades [1, 2, 3, 4, 5, 6, 7], breakthroughs in practical quantum algorithms remain constrained by prohibitive circuit depths [8, 9, 10, 11, 12, 13]. Although single-qubit and two-qubit gates form a universal set [14], they require substantially deeper circuits than direct synthesis of native multi-qubit gates [15, 16, 17]. The Controlled-Controlled-Z $(\text{CCZ})$ gate, a pivotal three-qubit operation, exemplifies this trade-off through its decomposition into elementary gates, which introduces substantial depth overhead and non-negligible fidelity degradation relative to direct implementation. This gate imparts a $\pi$ phase shift to targeted computational basis states—a property critical for Quantum Approximate Optimization Algorithm (QAOA[18], Grover’s search [16, 19, 16], and quantum error-correcting codes [20, 21, 22]. Consequently, realizing high-fidelity native CCZ gates on superconducting processors constitutes a pivotal challenge for advancing practical quantum computing.
@@
Figure 1: (a) Toffoli gate implementations: Conventional decomposition into single-qubit and six CZ gates (left) vs. direct implementation via a native three-qubit gate (right), achieving significant gate-count reduction. (b) Schematic of the superconducting quantum circuit: Three transmon qubits coupled via two tunable couplers. Each qubit is equipped with dedicated XY and Z control lines, each coupler with an independent Z control line, and individual readout resonators facilitate simultaneous state measurement. (c) Energy-level diagram of the three-qubit system after Schrieffer-Wolff transformation (SWT)
@@ II.2 Physical Model
(1) Three-qubit geometric phase accumulation: A genuine three-qubits entanglement operation is realized via nonadiabatic transitions between $|101\rangle$ and $|020\rangle$ states during the first stage. A key feature is the cyclic evolution back to the computational basis state $|101\rangle$, which accumulates a geometric $\pm\pi$ phase through a closed trajectory in Hilbert space.
@@
Figure 4: Calibration procedure. (a)(c) Calibration of transverse qubit-qubit coupling strength versus flux-bias voltage (V) in the coupler Z line. Top: Circuit schematic illustrating control lines and coupling elements. Bottom: $|101\rangle$ state population versus flux-bias voltage and time $\tau$, from which we extract the tansverse coupling strength. The dashed line marks the zero-coupling operating point. (b)(d) Calibration of the three-qubit CCZ gate. Scanning duration and amplitude of the Z pulse on $\mathrm{Qubit}2$, we measure the population of $|101\rangle$. And we identify th optimal operating point (marked as dashed line) where geometric phase accumulation reaches $\pi$. (e) Evolution trajectories of $|101\rangle$ and $|020\rangle$ states extracted from (d).
@@ III.1 Three-qubit geometric phase accumulation
For clarity, we hereafter denote a photon at frequency $\omega_{2}+\alpha_{2}$ by $\omega_{2}^{\prime}$, and express the joint state of $\mathrm{Qubit}\ 1,\mathrm{Qubit}\ 2,\mathrm{Qubit}\ 3$ by $|\mathrm{Qubit}\ 1,\mathrm{Qubit}\ 2,\mathrm{Qubit}\ 3\rangle$, where each qubit is described by its three lowest energy levels: the ground state $|0\rangle$, the first excited state $|1\rangle$ and the second excited state $|2\rangle$. Under the resonant condition $\omega_{1}+\omega_{3}=\omega_{2}+\omega_{2}^{\prime}$, a coherent three-body interaction emerges, enabling deterministic entanglement among the three qubits [32]. But in conventional settings, random arrival times of $\omega_{2}$ and $\omega_{2}^{\prime}$ would lead to destructive interference between the $U_{1}$ and $U_{2}$ processes within second-order perturbation (see Fig. 3(a)). As a result, the net two-photon transition matrix element vanishes [55].
To circumvent this interference, we employ cascaded transitions in superconducting artificial atoms to generate time–frequency entangled photon pairs. This approach ensures that $\omega_{2}^{\prime}$ arrives prior to $\omega_{2}$, thereby suppressing the $U_{2}$ process. Therefore, destructive interference is circumvented and three-body entanglement is deterministically generated. Based on this controlled interaction, we can achieve a high-fidelity three-qubit CCZ gate and proceed to compute the effective coupling strength $J$ of this two-photon transition [45], with $J$ being given by
@@
with $k_{ij}$ quantifying the coupling strength between $\omega_{i}$ and $\omega_{j}$. The condition $k_{ij}\ll|\omega_{i}-\omega_{j}|$ is naturally fulfilled in superconducting quantum circuits, strongly suppresses single-photon transitions and ensures the predominance of two-photon processes in the system’s evolution. This allows the derivation of an effective Hamiltonian for the three-qubit system:
@@ III.2 Active dynamical phase cancellation
Targeted CPhase gates are applied to cancel the residual ZZ and ZZZ couplings (as will be detailed in the Appendix. A). Through Ramsey-type interferometry measurements, we characterized and quantified the unwanted dynamical phases. Utilizing these target phase values, we then actively cancel them by applying non-adiabatic controlled-phase (CPhase) gates. Simulation results indicate that the required gate durations for these two cancellation pulses are $\tau_{12}$=23.4ns,and $\tau_{23}$=23.5ns respectively. We perform precision calibrations to enforce $\varphi_{13}+\varphi_{\mathrm{geometric}}=\pm\pi$. Given that $\varphi_{13}$ is a sufficiently small perturbation term in experimental realizations, we derive the approximation:
@@ V.2 Multi(five and beyond)-qubit Gate Extension
In conclusion, we propose a scheme to directly implement a CCZ gate on a superconducting circuit, achieving a simulated fidelity exceeding 99% within a gate time of 194 ns. The approach demonstrates exceptional robustness against experimental parameter drift and substantially reduced decoherence sensitivity, owing to its ultrashort duration that remains well below typical coherence times. Furthermore, the protocol incorporates a streamlined two-step calibration process that significantly reduces the deployment complexity on physical quantum processors, facilitating experimental realization and scalability. However, due to constraints of our available hardware setup, the experimental implementation of the CCZ gate is not realized in this study, forming a key objective for our subsequent work.
@@ Appendix A Calibration Procedures and Numerical Validation
We now proceed to calibrate the optimal parameters for the three-qubit CCZ interaction. This is achieved by fixing the frequencies of $\mathrm{Qubit}\ 1$ and $\mathrm{Qubit}\ 3$, along with the voltages applied to the $\mathrm{Coupler}\ 1$ and $\mathrm{Coupler}\ 2$, while scanning both the frequency of $\mathrm{Qubit}\ 2$ and the interaction time. The system’s state is subsequently measured to identify the optimal operational regime. The optimal operational point in this step can be readily identified from the two-dimensional scan—simply locating the star-marked point in Fig. 4(d) of the main text. A more critical task, however, is the accurate characterization of the dynamical phases accumulated during this process, which must be actively canceled in the subsequent calibration stage. As shown in a previous study [48], these phases can be measured through initializing three qubits in six special states, i.e., $|0\!+\!0\rangle$$|1\!+\!1\rangle$$\rangle|0+1|0\!+\!1\rangle$|1 0+|1\,0\ $\rangle$, and |0 0+|0\,0\ $\rangle$. We then measure all conditional phases through Ramsy-Like experiments, i.e., $\varphi_{12}=\varphi_{|1+0\rangle}-\varphi_{|0+0\rangle}$, $\varphi_{23}=\varphi_{|0+1\rangle}-\varphi_{|0+0\rangle}$, $\varphi_{13}=\varphi_{|1\,0\,\!+\!\rangle}-\varphi_{|0\,0\,\!+\!\rangle}$, $\varphi_{123}=\varphi_{|1+1\rangle}-\varphi_{|0+0\rangle}$, where $\varphi_{12},\varphi_{23},\varphi_{123}$ are defined in the main text. With these characterized parameters, we can eliminate the residual phases via targeted CPhase gate operations, thereby obtaining a purified CCZ gate.

### v2

# revtex4-2Repair the float

@@ I Introduction
Despite significant advances in superconducting quantum computing over the past few decades [1, 2, 3, 4, 5, 6, 7], breakthroughs in practical quantum algorithms remain constrained by prohibitive circuit depths [8, 9, 10, 11, 12, 13]. Although single-qubit and two-qubit gates form a universal set [14], they require substantially deeper circuits than direct synthesis of native multi-qubit gates [15, 16, 17]. The Controlled-Controlled-Z $(\text{CCZ})$ gate, a pivotal three-qubit operation, exemplifies this trade-off through its decomposition into elementary gates, which introduces substantial depth overhead and non-negligible fidelity degradation relative to direct implementation. This gate imparts a $\pi$ phase shift to targeted computational basis states—a property critical for Quantum Approximate Optimization Algorithm (QAOA[18], Grover’s search [16, 19, 16], and quantum error-correcting codes [20, 21, 22]. Consequently, realizing high-fidelity native CCZ gates on superconducting processors is critical for advancing practical quantum computing.
@@
Figure 1: (a) Toffoli gate implementations: Conventional decomposition into single-qubit and six CZ gates (left) vs. direct implementation via a native three-qubit gate (right), achieving significant gate-count reduction. (b) Schematic of the superconducting quantum circuit: Three transmon qubits coupled via two tunable couplers. Each qubit is equipped with dedicated XY and Z control lines, each coupler with an independent Z control line, and individual readout resonators facilitate simultaneous state measurement. (c) Energy-level diagram of the three-qubit system after Schrieffer-Wolff transformation (SWT).
@@ II.2 Physical Model
(1) Three-qubit geometric phase accumulation: A genuine three-qubits entanglement operation is realized via nonadiabatic transitions between $|101\rangle$ and $|020\rangle$ states during the first stage. A key feature is the cyclic evolution back to the computational basis state $|101\rangle$, which accumulates a geometric $\pi$ phase through a closed trajectory in Hilbert space.
@@
Figure 4: Calibration procedure. (a)(c) Calibration of transverse qubit-qubit coupling strength versus flux-bias voltage (V) in the coupler Z line. Top: Circuit schematic illustrating control lines and coupling elements. Bottom: $|101\rangle$ state population versus flux-bias voltage and time $\tau$, from which we extract the tansverse coupling strength. The dashed line marks the zero-coupling operating point. (b)(d) Calibration of the three-qubit CCZ gate. Scanning duration and amplitude of the Z pulse on $\mathrm{Qubit}2$, we measure the population of $|101\rangle$. And we identify th optimal operating point (marked by a red pentagram on the dashed line) where geometric phase accumulation reaches $\pi$. (e) Evolution trajectories of $|101\rangle$ and $|020\rangle$ states extracted from (d).
@@ III.1 Three-qubit geometric phase accumulation
For clarity, we hereafter denote a photon at frequency $\omega_{2}+\alpha_{2}$ by $\omega_{2}^{\prime}$, and express the joint state of $\mathrm{Qubit}\ 1,\mathrm{Qubit}\ 2,\mathrm{Qubit}\ 3$ by $|\mathrm{Qubit}\ 1,\mathrm{Qubit}\ 2,\mathrm{Qubit}\ 3\rangle$, where each qubit is described by its three lowest energy levels: the ground state $|0\rangle$, the first excited state $|1\rangle$ and the second excited state $|2\rangle$. Under the resonant condition $\omega_{1}+\omega_{3}=\omega_{2}+\omega_{2}^{\prime}$, a coherent three-body interaction emerges, enabling deterministic entanglement among the three qubits [32]. But in conventional settings, random arrival times of $\omega_{2}$ and $\omega_{2}^{\prime}$ would lead to destructive interference between the $U_{1}$ and $U_{2}$ processes within second-order perturbation (see Fig. 2(b)). As a result, the net two-photon transition matrix element vanishes [55].
To circumvent this interference, we employ cascaded transitions in superconducting artificial atoms to generate time–frequency entangled photon pairs. In Fig. 3, we illustrate the physical realization of this process using superconducting qubits, which is achieved by resonantly coupling the $|101\rangle$ and $|020\rangle$ states. This facilitates an energy exchange in which Qubit 2 loses two photons while Qubit 1 and Qubit 3 each absorb one, even though each absorption process is individually off-resonance. By facilitating this, the approach ensures that $\omega_{2}^{\prime}$ arrives prior to $\omega_{2}$, thereby suppressing the $U_{2}$ process. Therefore, destructive interference is circumvented and three-body entanglement is deterministically generated. Based on this controlled interaction, we can achieve a high-fidelity three-qubit CCZ gate and proceed to compute the effective coupling strength $J$ of this two-photon transition [45], with $J$ being given by
@@
where $k_{ij}$ quantifying the coupling strength between $\omega_{i}$ and $\omega_{j}$. The condition $k_{ij}\ll|\omega_{i}-\omega_{j}|$ is naturally fulfilled in superconducting quantum circuits, strongly suppresses single-photon transitions and ensures the predominance of two-photon processes in the system’s evolution. This allows the derivation of an effective Hamiltonian for the three-qubit system:
@@ III.2 Active dynamical phase cancellation
Targeted CPhase gates are applied to cancel the residual ZZ and ZZZ couplings (as will be detailed in the Appendix. A). Through Ramsey-type interferometry measurements, we characterized and quantified the unwanted dynamical phases. Utilizing these target phase values, we then actively cancel them by applying non-adiabatic controlled-phase (CPhase) gates. Simulation results indicate that the required gate durations for these two cancellation pulses are $\tau_{12}$ = 23.4 ns, and $\tau_{23}$ = 23.5 ns respectively. We perform precision calibrations to enforce $\varphi_{13}+\varphi_{\mathrm{geometric}}=\pi$. Given that $\varphi_{13}$ is a sufficiently small perturbation term in experimental realizations, we derive the approximation:
@@ V.2 Multi(five and beyond)-qubit Gate Extension
In conclusion, we propose a scheme to directly implement a CCZ gate on a superconducting circuit, achieving a simulated fidelity exceeding 99% within a gate time of 194 ns. The approach demonstrates exceptional robustness against experimental parameter drift and substantially reduced decoherence sensitivity, owing to its short duration that remains well below typical coherence times. Furthermore, the protocol incorporates a streamlined two-step calibration process that significantly reduces the deployment complexity on physical quantum processors, facilitating experimental realization and scalability. However, due to constraints of our available hardware setup, the experimental implementation of the CCZ gate is not realized in this study, forming a key objective for our subsequent work.
@@ Appendix A Calibration Procedures and Numerical Validation
Subsequently, we describe the procedure for determining the optimal parameters required to realize the three-qubit CCZ interaction. This is achieved by fixing the frequencies of $\mathrm{Qubit}\ 1$ and $\mathrm{Qubit}\ 3$, along with the voltages applied to the $\mathrm{Coupler}\ 1$ and $\mathrm{Coupler}\ 2$, while scanning both the frequency of $\mathrm{Qubit}\ 2$ and the interaction time. The system’s state is subsequently measured to identify the optimal operational regime. The optimal operational point in this step can be readily identified from the two-dimensional scan—simply locating the star-marked point in Fig. 4(d) of the main text. A more critical task, however, is the accurate characterization of the dynamical phases accumulated during this process, which must be actively canceled in the subsequent calibration stage. As shown in a previous study [48], these phases can be measured through initializing three qubits in six special states, i.e., $|0\!+\!0\rangle$$|1\!+\!1\rangle$$\rangle|0+1|0\!+\!1\rangle$|1 0+|1\,0\ $\rangle$, and |0 0+|0\,0\ $\rangle$. We then measure all conditional phases through Ramsy-Like experiments, i.e., $\varphi_{12}=\varphi_{|1+0\rangle}-\varphi_{|0+0\rangle}$, $\varphi_{23}=\varphi_{|0+1\rangle}-\varphi_{|0+0\rangle}$, $\varphi_{13}=\varphi_{|1\,0\,\!+\!\rangle}-\varphi_{|0\,0\,\!+\!\rangle}$, $\varphi_{123}=\varphi_{|1+1\rangle}-\varphi_{|0+0\rangle}$, where $\varphi_{12},\varphi_{23},\varphi_{123}$ are defined in the main text. With these characterized parameters, we can eliminate the residual phases via targeted CPhase gate operations, thereby obtaining a purified CCZ gate.